The perimeter of a rectangle is 30m. The length of the rectangle is 3m more than twice the width. Find the dimensions of the rectangle.
Answer by richard1234(7193)

Suppose the width is w, and the length is 2w + 3. We now have:
2(w + 2w + 3) = 30
3w + 3 = 15
w = 4
L = 2w + 3 = 11.